# Pareto Envelope Augmented with Reinforcement Learning: Multi-objective reinforcement learning-based approach for

### 19 Dec 2023 v2

@@ Abstract
A novel method, the Pareto Envelope Augmented with Reinforcement Learning (PEARL), has been developed to address the challenges posed by multi-objective problems, particularly in the field of engineering where the evaluation of candidate solutions can be time-consuming. PEARL distinguishes itself from traditional policy-based multi-objective Reinforcement Learning methods by learning a single policy, eliminating the need for multiple neural networks to independently solve simpler sub-problems. Several versions inspired from deep learning and evolutionary techniques have been crafted, catering to both unconstrained and constrained problem domains. Curriculum Learning (CL) is harnessed to effectively manage constraints in these versions. PEARL’s performance is first evaluated on classical multi-objective benchmarks. Additionally, it is tested on two practical PWR core Loading Pattern (LP) optimization problems to showcase its real-world applicability. The first problem involves optimizing the Cycle length ($L_{C}$) and the rod-integrated peaking factor ($F_{\Delta h}$) as the primary objectives, while the second problem incorporates the mean average enrichment as an additional objective. Furthermore, PEARL addresses three types of constraints related to boron concentration ($C_{b}$), peak pin burnup ($Bu_{max}$), and peak pin power ($F_{q}$). The results are systematically compared against a conventional approach, the Non-dominated Sorting Genetic Algorithm (NSGA). Notably, PEARL, specifically the PEARL-NdS (crowding) variant, efficiently uncovers a Pareto front without necessitating additional efforts from the algorithm designer, as opposed to a single optimization with scaled objectives. It also outperforms the classical approach across multiple performance metrics, including the Hyper-volume. Future works will encompass a sensitivity analysis of hyper-parameters with statistical analysis to optimize the application of PEARL and extend it to more intricate problems.
@@
$Bu_{max}$ Peak pin burnup $C_{b}$ Boron concentration $F_{q}$ Peak pin power $F_{\Delta h}$ Rod integrated peaking factor $L_{C}$ Cycle Length CL Curriculum Learning CO Combinatorial Optimization EOC End Of Cycle ET Evolutionary Techniques FA Fuel Assembly FOMs Figure Of Merits HV Hyper-Volume LP Loading Pattern MDP Markov Decision Process MOO Multi-Objective Optimization NSGA Non-dominated Sorting Genetic Algorithm PEARL Pareto Envelope Augmented with Reinforcement Learning PPO Proximal Policy Optimization PWR Pressurized Water Reactor RL Reinforcement Learning SO Stochastic Optimization
@@ 1 Introduction
Deep learning and in particular deep reinforcement learning (DRL) [8] are frameworks that do not work with analytical representation of an objective function but with objects (e.g. function approximator) that are constructed statistically with a large number of data and then mathematically optimized [9]. There are two classical ways to leverage deep learning for optimization: either replacing the function evaluation (e.g., a simulator) by a surrogate, or use it to learn a set of decision rules about the underlying structure of a problem as for RL to generalize from seen to unseen instances of the problems [9, 1]. In the context of large-scale CO, a policy (or sequence of decision) is approximated as a neural network, which weights are trained to generate solution end-to-end of unseen instances of problems such as the Traveling Salesman (TSP) or the Knapsack (KP) very fast [10]. A simple forward pass from the trained networks that implement $P(Y|X)=\prod_{k=1}^{N}P(y_{k+1}|y_{1},..,y_{k},X_{k})$, where $X=(x_{1},...,x_{N})\in\mathbb{R}^{p}$ is an input embedding and $Y=(y_{1},...,y_{N})$ is a permutation of elements, provides the solution. With the concept of efficient generalization in mind, Deep RL (DRL) has demonstrated remarkable prowess in handling classical CO problems, allowing it to generalize effectively from observed instances to previously unseen ones [11]. This feature has also found valuable application in MOO alongside RL. For instance, in [12], the TSP is decomposed into N (= 100) scalar objectives obtained from the weighted sum of the different tour costs, each of which is modeled as its neural network. A Neighborhood-based parameter-transfer strategy is utilized to uncover the Pareto front: all sub-problems are solved sequentially and the optimized weights of each neural network are transferred to the next, as neighboring points on the Pareto surface should be close enough, which accelerates the search of new points on the surface. The authors of this study assert that their method surpasses classical heuristic-based MOO methods in terms of computation time, the number of solutions found, and Hyper-Volume (HV). Further instances of studies and techniques for applying MOO with RL, as well as their limitations in the context of our specific problem, are elaborated upon in Section 2.3. Notably, a critical limitation, which was emphasized in our previous study [2], arises when implementing an end-to-end inference strategy for MOO, particularly when the objective function demands substantial computational time for evaluation [13, 2]. Similar to the single-objective scenario, there was a pressing need to identify alternative strategies for harnessing RL’s potential in MOO, specifically within the realm of engineering design.
@@
The effectiveness and applicability of PEARL are rigorously assessed through a series of comprehensive tests on established classical multi-objective benchmark problems: the dtlz test suite [21] for the unconstrained cases, and the cxdtlzy [17] and ctpx [22] test suite for the constrained cases. Then, we compare their performance on a typical Pressurized Water Reactor (PWR) design [1, 2] against classical SO-based approaches: NSGA-II [15] and -III [16, 17], which are established approaches used in Genetic Algorithm (GA)-based multi-objective optimization [3, 4]. These methods have a proven track record in various applications within the nuclear field, including neutron spectrum estimation and advanced reactor optimization [3]. A summary of the algorithm developed and tested for this work for the unconstrained and constrained cases is given in table 1:
Table 1: Summary of the algorithms developed and tested in this work.
@@
Unconstrained cos,KL-div $\epsilon$ crowding,niching
Constrained -CL -CL ${}^{\star}$,-niching2
@@
    Pioneering application of multi-objective RL, both in unconstrained and constrained settings, to the core LP problem. This extends the groundwork laid in our previous work on single-objective deep RL for the same problem [1, 18].
@@ 3.1 Unconstrained PEARL
where $\tilde{r_{i}(s_{t},a_{t})}=\frac{w_{j}r_{i}}{\sum_{k}w_{k}r_{k}}$. $\mu_{u}$ estimates the extent of which $\tilde{r}$ is non-uniform. Alternatively, in [31], the factor $\mu_{u}$ is equal to $-\lambda\frac{w^{T}J(x,y,r;\theta)}{||w||\cdot||J(x,y,r;\theta)||}$ (i.e., the cosine similarity), where $\lambda$ is a weighting coefficient to tune.
@@
Our second set of methods is inspired by SO and in particular Evolutionary Techniques (ET). We will leverage ET formulations to shape the reward which serves as a measure of the quality of the information provided by a newly generated solution on the existing approximation of the Pareto front. One such reward can be indicator-based. The additive $\epsilon-$indicator $I_{\epsilon}^{+}$ (e.g., IBEA [28]) is the method of choice due to its popularity [25] and the fact that it is inexpensive to compute [28]. Let a population of known pareto equivalent solution $\mathcal{P}$ and $F_{I_{\epsilon}^{+}}(x_{t})$ the fitness of the new solution generated $x_{t}$ at time step t. $F_{I_{\epsilon}^{+}}(x_{t})$ represents the loss of quality by adding $x_{t}$ to the population (see [28] for the exact definition). In our case, the population is a buffer of non-dominated solutions encountered during the search because, unlike the original ET-based paper [16], we do not have a population to play the role of the optimum front. Our choice comes from MOSA [6], where a reference buffer was used to overcome this issue. First, any solution $(y_{i},r_{i})_{i\in\{1,...,k\}}$ dominated by the new solution $(x_{t},r_{t})$ is removed, while if $x_{t}$ is dominated by any of the existing solutions, it is not kept and assigned a negative reward equal to the maximum size of the buffer of Pareto-equivalent solutions $\kappa$ specified by the user (equivalent to the population size in IBEA). Otherwise, we measure $F_{I_{\epsilon}^{+}}(y_{i})$ $\forall y_{i}\in\mathcal{P}$ and rank $x$ based on $F_{I_{\epsilon}^{+}}(x_{t})$. This method is labeled PEARL-$\epsilon$. Alternatively, we could assign a reward based on the distribution of solution on the pareto front, which draw inspiration from the density-based methods in multi-objective GA, namely NSGA-II [15] and NSGA-III [16, 17]. Therefore, we modified the NSGA-II/III ranking assignment by also applying a density-based ranking procedure to a buffer of Pareto equivalent solutions and a new solution $x$. This method is labeled PEARL-Nondominated Sorting (PEARL-NdS). When the method is based on NSGA-II and NSGA-III, we note PEARL-NdS (crowding) and PEARL-NdS (niching), respectively.
@@ 4.1 Tools for performance comparisons
Measuring the performance of a multi-objective algorithm presents a more intricate challenge compared to evaluating a single-objective one. This complexity arises because a single scalar value cannot provide a comprehensive quantitative assessment of the algorithm’s quality. In the literature, the four most utilized are the HV, the $\epsilon$-indicator ($I_{\epsilon}^{+}$), the Generational Distance (GD), and the Inverted Generational Distance (IGD) [25]. The HV is a unary metric that only requires a reference point that should be dominated by all points in $\mathcal{P}^{\star}$, which is the nadir in [5]. A higher HV is better. The other three are binary metrics that require the knowledge of all the elements of the approximate Pareto-optimal set (although some alternatives exist [25]). For the latter three a lower value is better. In accordance with existing literature, we utilize the Pareto front generated by running all algorithms as the reference set of solutions [5, 26]. We will also consider the C_metric(A,Z) and $I_{C}$, which is the average number of point in set A dominated by points in the reference set Z and the number of them, respectively. A higher value the better. Furthermore, it’s important to note that not all Pareto-optimal points are of equal importance [7]. Even if we have a set of Pareto-optimal solutions, we could propose a scheme to offer the best solution for the decision-maker. We will use the entropy method as described in [7].
In the classical benchmark study presented in Section 5.1, we provide the HV and cardinality of Pareto-optimal solutions (via the $I_{C}$). However, we do not provide the $I_{\epsilon}^{+}$, neither GD, IGD, nor C_metric in this context because the Pareto front is known, generated automatically, and depends on the desired number of points. The primary objective is to determine whether the points generated by our algorithm are dominated by these known solutions. On the other hand, for the PWR benchmark discussed in Section 5.2, we will utilize all the mentioned metrics to comprehensively assess algorithm performance.
@@ 4.2.1 Classical benchmarks
To demonstrate the validity and performance of the developed algorithms, we will subject them to testing across various multi-objective benchmarks, both in unconstrained and constrained settings. For the unconstrained benchmarks, our selection includes the dtlzX test suite [21], specifically dtlz2, dtlz4, dtlz5, dtlz6, and dtlz7. dtlz2 presents a dense Pareto front, while dtlz4 assesses an algorithm’s ability to maintain a well-distributed set of solutions. dtlz5 exhibits an optimal front in the form of a curve, and dtlz6 introduces added complexity with a non-linear distance function constraint g, making convergence against the Pareto-optimal curve more challenging. dtlz7 features disconnected Pareto-optimal regions, thereby evaluating the algorithm’s capability to maintain sub-populations across different Pareto-optimal regions [21].
In terms of constrained benchmarks, we opted to utilize classical benchmarks namely the cxdtlzy [17] and ctpx [22] suites, which detailed can be found on pymoo’s website [47]. To vary the shape and type of constrained spaces, we chose to study c2-dtlz2 (type 2 in [17]) and c3-dtlz4 (type 3 in [17]). The ctpx test suite [22] is also interesting. In ctp1, the infeasible region is very near the Pareto front, ctp2 is an ensemble of discrete Pareto regions, ctp3 reduces the feasible Pareto region to a single point, while ctp4 increases the difficulty of ctp3 by making the transition from continuous to discontinuous feasible spaces further away from the Pareto-optimal points [22].
Fortunately for these cases, we have access to the optimal Pareto front. Therefore, to assess the performance of the PEARL algorithms, we will utilize the Pareto front available in pymoo [47].
@@ 4.2.3 Legacy Approaches used for comparison
To compare every version of PEARL (i.e., PEARL-e, PEARL-$\epsilon$, and PEARL-NdS) against legacy approaches for the unconstrained case and C-PEARL-NdS with the formulation in equation 8 for the $CV$ in the constrained case, we adopted NSGA-II [15] (which constrained version was not developed) and NSGA-III [16, 17]. They are often considered the state-of-the-art GA-based approach for multi-objective optimization [3]. We also use PEARL-NdS (crowding2) and PEARL-NdS (niching2) to compare the performance of a pure ranking versus a smoother satisfaction of the constraints with $CV$ as a penalty in PEARL-NdS (crowding) and PEARL-NdS (niching).

### 15 Mar 2024 v3

# Multi-Objective Reinforcement Learning-based Approach for Pressurized Water Reactor Optimization

@@ Abstract
A novel method, the Pareto Envelope Augmented with Reinforcement Learning (PEARL), has been developed to address the challenges posed by multi-objective problems, particularly in the field of engineering where the evaluation of candidate solutions can be time-consuming. PEARL distinguishes itself from traditional policy-based multi-objective Reinforcement Learning methods by learning a single policy, eliminating the need for multiple neural networks to independently solve simpler sub-problems. Several versions inspired from deep learning and evolutionary techniques have been crafted, catering to both unconstrained and constrained problem domains. Curriculum Learning (CL) is harnessed to effectively manage constraints in these versions. PEARL’s performance is first evaluated on classical multi-objective benchmarks. Additionally, it is tested on two practical PWR core Loading Pattern (LP) optimization problems to showcase its real-world applicability. The first problem involves optimizing the Cycle length ($L_{C}$) and the rod-integrated peaking factor ($F_{\Delta h}$) as the primary objectives, while the second problem incorporates the mean average enrichment as an additional objective. Furthermore, PEARL addresses three types of constraints related to boron concentration ($C_{b}$), peak pin burnup ($Bu_{max}$), and peak pin power ($F_{q}$). The results are systematically compared against conventional approaches from stochastic optimization. Notably, PEARL, specifically the PEAL-NdS (crowding) variant, efficiently uncovers a Pareto front without necessitating additional efforts from the algorithm designer, as opposed to a single optimization with scaled objectives. It also outperforms the classical approach across multiple performance metrics, including the Hyper-volume. Future works will encompass a sensitivity analysis of hyper-parameters with statistical analysis to optimize the application of PEARL and extend it to more intricate problems.
@@
$Bu_{max}$ Peak pin burnup $C_{b}$ Boron concentration $F_{q}$ Peak pin power $F_{\Delta h}$ Rod integrated peaking factor $L_{C}$ Cycle Length CL Curriculum Learning CO Combinatorial Optimization EOC End Of Cycle ET Evolutionary Techniques FA Fuel Assembly FOMs Figure Of Merits HV Hyper-Volume LP Loading Pattern MDP Markov Decision Process MOO Multi-Objective Optimization NSGA Non-dominated Sorting Genetic Algorithm PEARL Pareto Envelope Augmented with Reinforcement Learning PPO Proximal Policy Optimization PWR Pressurized Water Reactor RL Reinforcement Learning SA Simulated Annealing SO Stochastic Optimization TS Tabu Search
@@ 1 Introduction
Deep learning and in particular deep reinforcement learning (DRL) [8] are frameworks that do not work with analytical representation of an objective function but with objects (e.g. function approximator) that are constructed statistically with a large number of data and then mathematically optimized [9]. There are two classical ways to leverage deep learning for optimization: either replacing the function evaluation (e.g., a simulator) by a surrogate, or use it to learn a set of decision rules about the underlying structure of a problem as for RL to generalize from seen to unseen instances of the problems [1, 2, 9]. In the context of large-scale CO, a policy (or sequence of decision) is approximated as a neural network, which weights are trained to generate solution end-to-end of unseen instances of problems such as the Traveling Salesman (TSP) or the Knapsack (KP) very fast [10]. A simple forward pass from the trained networks that implement $P(Y|X)=\prod_{k=1}^{N}P(y_{k+1}|y_{1},..,y_{k},X_{k})$, where $X=(x_{1},...,x_{N})\in\mathbb{R}^{p}$ is an input embedding and $Y=(y_{1},...,y_{N})$ is a permutation of elements, provides the solution. With the concept of efficient generalization in mind, Deep RL (DRL) has demonstrated remarkable prowess in handling classical CO problems, allowing it to generalize effectively from observed instances to previously unseen ones [11]. This feature has also found valuable application in MOO alongside RL. For instance, in [12], the TSP is decomposed into N (= 100) scalar objectives obtained from the weighted sum of the different tour costs, each of which is modeled as its neural network. A Neighborhood-based parameter-transfer strategy is utilized to uncover the Pareto front: all sub-problems are solved sequentially and the optimized weights of each neural network are transferred to the next, as neighboring points on the Pareto surface should be close enough, which accelerates the search of new points on the surface. The authors of this study assert that their method surpasses classical heuristic-based MOO methods in terms of computation time, the number of solutions found, and Hyper-Volume (HV). Further instances of studies and techniques for applying MOO with RL, as well as their limitations in the context of our specific problem, are elaborated upon in Section 2.3. Notably, a critical limitation, which was emphasized in our previous study [2], arises when implementing an end-to-end inference strategy for MOO, particularly when the objective function demands substantial computational time for evaluation [13, 2]. Similar to the single-objective scenario, there was a pressing need to identify alternative strategies for harnessing RL’s potential in MOO, specifically within the realm of engineering design.
@@
The effectiveness and applicability of PEARL are rigorously assessed through a series of comprehensive tests on established classical multi-objective benchmark problems: the dtlzX test suite [21] for the unconstrained cases, and the cxdtlzy [17] and ctpx [22] test suite for the constrained cases. Then, we compare their performance on a typical Pressurized Water Reactor (PWR) design [1, 2] against classical SO-based approaches: NSGA-II [15] and -III [16, 17], Simulated Annealing (SA), and Tabu Search (TS), which are established approaches used for multi-objective optimization[3, 4, 5, 6]. These methods have a proven track record in various applications within the nuclear field, including neutron spectrum estimation and advanced reactor optimization [3, 23]. A summary of the RL-based algorithms developed and tested for this work for the unconstrained and constrained cases is given in table 1:
Table 1: Summary of the algorithms developed and tested in this work. ’-CL’ in the constrained case means that the constraints are first solved in a single-objective manner and then only the ,ulti-objective version is called.
@@
Unconstrained variant(s) cos,KL-div $\epsilon$ crowding,niching
Constrained variant(s) -CL -CL ${}^{\star}$,-niching2
@@
    Pioneering application of multi-objective RL, both in unconstrained and constrained settings, to the core LP problem. This extends the groundwork laid in our previous work on single-objective deep RL for the same problem [1, 2, 18].
@@ 3.1 Unconstrained PEARL
[…]
@@
Our second set of methods is inspired by SO and in particular Evolutionary Techniques (ET). We will leverage ET formulations to shape the reward which serves as a measure of the quality of the information provided by a newly generated solution on the existing approximation of the Pareto front. One such reward can be indicator-based. The additive $\epsilon-$indicator $I_{\epsilon}^{+}$ (e.g., IBEA [29]) is the method of choice due to its popularity [26] and the fact that it is inexpensive to compute [29]. Let a population of known pareto equivalent solution $\mathcal{P}$ and $F_{I_{\epsilon}^{+}}(x_{t})$ the fitness of the new solution generated $x_{t}$ at time step t. $F_{I_{\epsilon}^{+}}(x_{t})$ represents the loss of quality by adding $x_{t}$ to the population (see [29] for the exact definition). In our case, the population is a buffer of non-dominated solutions encountered during the search because, unlike the original ET-based paper [16], we do not have a population to play the role of the optimum front. Our choice draws inspiration from MOSA [6] and or Pareto Evelope-based Selection Algorithm (PESA) [40], where a reference buffer was used to overcome this issue. First, any solution $(y_{i},r_{i})_{i\in\{1,...,k\}}$ dominated by the new solution $(x_{t},r_{t})$ is removed, while if $x_{t}$ is dominated by any of the existing solutions, it is not kept and assigned a negative reward equal to the maximum size of the buffer of Pareto-equivalent solutions $\kappa$ specified by the user (equivalent to the population size in IBEA). Otherwise, we measure $F_{I_{\epsilon}^{+}}(y_{i})$ $\forall y_{i}\in\mathcal{P}$ and rank $x$ based on $F_{I_{\epsilon}^{+}}(x_{t})$. This method is labeled PEARL-$\epsilon$. Alternatively, we could assign a reward based on the distribution of solution on the pareto front, which draw inspiration from the density-based methods in multi-objective GA, namely NSGA-II [15] and NSGA-III [16, 17]. Therefore, we modified the NSGA-II/III ranking assignment by also applying a density-based ranking procedure to a buffer of Pareto equivalent solutions and a new solution $x$. This method is labeled PEARL-Nondominated Sorting (PEARL-NdS). When the method is based on NSGA-II and NSGA-III, we note PEARL-NdS (crowding) and PEARL-NdS (niching), respectively.
@@ 4.1 Tools for performance comparisons
Measuring the performance of a multi-objective algorithm presents a more intricate challenge compared to evaluating a single-objective one. This complexity arises because a single scalar value cannot provide a comprehensive quantitative assessment of the algorithm’s quality. Because of this complexity, it is always better to utilize multiple ones to rigorously assess the performances of multi-objective approaches [49]. In the literature, the four most utilized are the HV, the $\epsilon$-indicator ($I_{\epsilon}^{+}$), the Generational Distance (GD), and the Inverted Generational Distance (IGD) [26]. The HV is a unary metric that only requires a reference point that should be dominated by all points in $\mathcal{P}^{\star}$, which is the nadir in [5]. A higher HV is better. The other three are binary metrics that require the knowledge of all the elements of the approximate Pareto-optimal set (although some alternatives exist [26]). For the latter three a lower value is better. Similarly, because the IGD is not Pareto-compliant, i.e., some set B dominated by another set A can have a better IGD value (i.e., lower value) than A [49], we have also added the ${}^{+}$. The ${}^{+}$ and provide better assessment than its IGD counterpart because it depends less on the distribution of the reference solution set [49] (i.e., it is weakly Pareto-compliant). In accordance with existing literature, we utilize the Pareto front generated by running all algorithms as the reference set of solutions [5, 27]. We will also consider the C_metric(A,Z) and $I_{C}$, which is the average number of point in set A dominated by points in the reference set Z and the number of them, respectively. A higher value the better. Furthermore, it’s important to note that not all Pareto-optimal points are of equal importance [7]. Even if we have a set of Pareto-optimal solutions, we could propose a scheme to offer the best solution for the decision-maker. We will use the entropy method as described in [7].
In the classical benchmark study presented in Section 5.1, we provide the HV and cardinality of Pareto-optimal solutions (via the $I_{C}$). However, we do not provide the $I_{\epsilon}^{+}$, neither GD, IGD, ${}^{+}$, nor C_metric in this context because the Pareto front is known, generated automatically, and depends on the desired number of points. The primary objective is to determine whether the points generated by our algorithm are dominated by these known solutions. On the other hand, for the PWR benchmark discussed in Section 5.2, we will utilize all the mentioned metrics to comprehensively assess algorithm performance.
@@ 4.2.1 Classical benchmarks
To demonstrate the validity and performance of the developed algorithms, we will subject them to testing across various multi-objective benchmarks, both in unconstrained and constrained settings. For the unconstrained benchmarks, our selection includes the dtlzX test suite [21], specifically dtlz1, dtlz2, dtlz3, dtlz4, dtlz5, dtlz6, and dtlz7. dtlz1 is an M-objective problem with linear Pareto-optimal front. dtlz2 presents a dense Pareto front, while dtlz3 assesses an algorithm’s ability to converge to a global pareto front with many local optima an MOO approach can converge to. dtlz4 assesses an algorithm’s ability to maintain a well-distributed set of solutions. dtlz5 exhibits an optimal front in the form of a curve, and dtlz6 introduces added complexity with a non-linear distance function constraint g, making convergence against the Pareto-optimal curve more challenging. dtlz7 features disconnected Pareto-optimal regions, thereby evaluating the algorithm’s capability to maintain sub-populations across different Pareto-optimal regions [21].
In terms of constrained benchmarks, we opted to utilize classical benchmarks namely the cxdtlzy [17] and ctpx [22] suites, which detailed can be found on pymoo’s website [50]. To vary the shape and type of constrained spaces, we chose to study c1dtlz1 and c1dtlz3 (type 1 in [17]), c2dtlz2 (type 2 in [17]), and c3dtlz4 (type 3 in [17]). The ctpx test suite [22] is also interesting. In ctp1, the infeasible region is very near the Pareto front, ctp2 is an ensemble of discrete Pareto regions, ctp3 reduces the feasible Pareto region to a single point, while ctp4 increases the difficulty of ctp3 by making the transition from continuous to discontinuous feasible spaces further away from the Pareto-optimal points [22].
Fortunately for these cases, we have access to the optimal Pareto front. Therefore, to assess the performance of the PEARL algorithms, we will utilize the Pareto front available in pymoo [50]. Following the taxonomy of pymoo, the dimension of the input space is $n_{x}=n_{obj}+k-1$, where $n_{obj}$ is the number of objective for a particular problem and k is arbitrary. Because they are hard problems to solve, we choose k equal to 1 for dtlz1, dtlz3, c1dtlz1, and c1dtlz3. For the other unconstrained and constrained test cases we have k equal to 10 and 5, respectively.
@@ 4.2.3 Legacy Approaches used for comparison
For the core loading pattern optimization problem, the most widely leveraged heuristic-based algorithms are Genetic Algorithm (GA), Simulated Annealing (SA), and Tabu Search (TS) [52, 53, 54] whether in single- and multi-objective settings. Therefore, to compare every version of PEARL (i.e., PEARL-e, PEARL-$\epsilon$, and PEARL-NdS) against legacy approaches for the unconstrained case and C-PEARL-NdS with the formulation in equation 8 for the $CV$ in the constrained case, we adopted their multi-objective extensions:
  1. 1.
    For GA we utilize NSGA-II [15] (which constrained version was not developed) and NSGA-III [16, 17]. They are often considered the state-of-the-art GA-based approach for multi-objective optimization [3] and are heavily leveraged in the field of nuclear science and engineering [23, 55].
  2. 2.
    For SA, we use the formulation of [6], where the acceptance of the new candidate hinge upon Pareto equivalence relationships between this solution and a buffer of stored solutions, as well as the $CV$. In case the previous $x$ and new candidates $\mathcal{N}(x)$ are both infeasible, the probability of acceptance is $\rho\sim e^{\frac{CV(x)-CV(\mathcal{N}(x))}{T}}$. In [6], the Temperate T is constant, but we adopted an adaptive formulation that changes T depending on the values of each constraint and objective as in other core loading pattern optimization works [52, 56].
  3. 3.
    For TS, we draw inspiration from recent work on unconstrained MOO with TS [5], which we extend to constrained optimization with [57]; The Ph.D. work in [5] is built on top of [57] and extended successfully the single-objective application of TS to core loading pattern optimization from [58] within the PANTHER framework. The advantage was to allow unrestricted application of any kind of physic code and algorithm to their core loading problem rather than limiting it to the use of Generalized Perturbation Theory (GPT) [59]. Nevertheless, this work did not consider parallel implementation, nor did it implement constraints. The parallel implementation is explicated in a concurrent paper [54] implementing parallel Tabu Search (TS) for NEORL [60]. Moreover, our constraints’ handling approach is adopted from [57]. There, they assumed that all non-feasible solutions are Tabu, meaning that solutions are discarded until a feasible solution is found. We anticipate that we would have difficulties finding a feasible solution for real-world problems by random sampling. Indeed, our experience with single-objective and unconstrained multi-objective demonstrated that downhill moves are favorable for future gains. Therefore, we implemented two options: One is the same as discussed above. For the other one, we utilize the $CV$ in a constrained non-domination criterion as described for NSGA-III and SA. In practice, we found that the second option may work better for the some continuous problems studied. Lastly, it is common for TS algorithms to have a Medium Term Memory (MTM) [5]. A MTM stores a certain number of high-quality moves that were seen but not taken (to the benefit of an even better one). We control this by the hyper-parameter intensification, which represents the probability of selecting a move from the MTM or the best current move.
To sum up, the multi-objective versions of SA and TS are similar to the single-objective ones, where the magnitude of the objective function is replaced by a comparison with a modified non-domination criterion as described in algorithm 3. This criterion is used also in NSGA but for selecting the parent population [17]:
Algorithm 3 Pseudocode for modified non-domination criterion for constrained multi-objective optimization
1:Input: E_prev, E
2:Ouput: E_prev $\succ$ E
3:if CV(E) $>$ 0 and CV(E_prev) == 0 then
4:    •return True
5:end if
6:if CV(E) == 0 and CV(E_prev) $>$then
7:    •Return False
8:end if
9:if CV(E) $>$ 0 and CV(E_prev) $>$then
10:    if SA then
11:         •Compute $\rho\sim e^{\frac{CV(E\_prev)-CV(E)}{T}}$
12:         •Compute $\epsilon\sim\mathcal{N}(0,1)$ return $\epsilon\leq\rho$
13:    else
14:         if CV(E) $>$ CV(E_prev) then
15:             •Return True
16:         else
17:             •Return False
18:         end if
19:    end if
20:end if
21:if CV(E) == 0 and CV(E_prev) == 0 then
22:    •use the classical unconstrained domination criteria
23:end if
Additionally, parallelism is implemented at the evaluation of each candidate solution in NSGA-II and NSGA-III. For both SA and TS, one processor is assigned to each Markov Chain [54]. There, each chain carries a local buffer corresponding to some high-quality solutions that new solutions generated must be compared with as in PEARL (see Section 3.1) or multi-objective SA [6] and PESA [40]. Nevertheless, we cannot restart each chain with one solution based on the magnitude of the best solutions found as in single-objective [54] but we must account for all solutions in the buffer. Therefore, at the end of the chain, the global Pareto front across all of them is first computed and a random sample is drawn to restart each.
We also defined PEARL-NdS (crowding2) and PEARL-NdS (niching2) to compare the performance of a pure ranking versus a smoother satisfaction of the constraints with $CV$ as a penalty in PEARL-NdS (crowding) and PEARL-NdS (niching). All of the algorithms described in this section as well as the variants of PEARL will be available in a further release of NEORL [60].